# Multi-Agent, Human-Agent and Beyond: A Survey on Cooperation in Social Dilemmas

### v1

@@ 4 Human-Agent Cooperation in SDs
This section explores human-agent cooperation in SDs from two perspectives: (1) developing algorithms that effectively enhance cooperative behavior; (2) the biases humans exhibit in their decision-making processes when interacting with both humans and agents.
@@ 4.1 Designing Algorithms for Cooperation with Humans
From S++ to S#.
Successful AI agents require three key properties: generality across various games, adaptability to unfamiliar counterparts, and swift learning in decision-making (Crandall et al., 2018). The S++ and S# family addresses these needs, focusing on iterated interactions where individuals repeatedly engage in the same game structure and make decisions based on past experiences. (Crandall, 2014; Oudah et al., 2018).
S++ (Crandall, 2014), an algorithm originally designed for iterated normal-form games, stands out for its two key features. First, it includes a range of experts, such as followers and model-based reinforcement learners, derived from the description of the game environment. Second, it employs an expert-selection mechanism, which determines the appropriate expert to be used in each game round. S++ outperforms other algorithms in interactions, including with model-based (mode-free) reinforcement learning algorithms, and exhibits as proficient at establishing cooperative relationships with humans as humans are (Crandall et al., 2018). However, high-level cooperation with humans often falls short, leading to the introduction of a communication mechanism.
S#, an advancement of S++, incorporates a communication framework enabling agents to engage in cheap talk, enhancing cooperation in simulations and human-agent experiments. The results demonstrate its ability to cooperate with humans and other algorithms at levels that rival human cooperation. Subsequently, Oudah et al. (2018) further develops the S++ and S# algorithms, introducing varied signaling strategies based on different philosophical approaches. Their experiments show that agents combining a behavioral strategy that learns quickly and effectively with a signaling strategy grounded in Carnegie’s Principles and explainable AI are most effective in winning friends and influencing people.
@@
Mitigating the obstacle of cooperation requires handling human selfishness. Zero-determinant strategy becomes a viable option. It includes extortion strategies, ensuring a player consistently outperforms its counterpart by a fixed percentage, and generosity strategies, aiming to incentivize cooperation from the counterpart. Hilbe et al. (2014) observe that although extortion strategies do prevail over their human counterparts, retaliation often follows, leading to defection. Consequently, generosity proves more profitable and effective. In contrast, extortion strategies excel in longer interactions, especially when humans are aware they interact with a computerized counterpart (Wang et al., 2016). In a safe control context, Zhang et al. (2023b) propose an algorithm that can prevent AI agents from being exploited by self-seeking humans and seeks to achieve a more optimal balance between safety and performance in human-agent interactions.
Experiment-driven insights.
Designing agents that promote cooperation can be inspired by experimental findings. Santos et al. (2020) initiate human-agent experiments and find that humans prefer cooperative partners after falling into collective failures, without expressing a significant preference for partners after collective successes. Expanding on this insight, they develop an evolutionary game theoretical model and validate the effectiveness of the partner selection rule. An online human-agent study reinforces these findings and highlights the significant influence of outcome-based strategies in human-agent interactions.
@@ 4.2 Revealing Human Bias towards Agents
Bias caused by the nature of counterparts.
In human-agent systems, revealing the true nature of agents can influence human behavior. Individuals tend to cooperate more with humans than agents (Karpus et al., 2021). Ishowo-Oloko et al. (2019) conduct an iterated PD experiment where humans are either accurately informed about the nature of their counterparts (humans or agents) or given false information. Employing the S++ algorithm, agents are more successful in eliciting human cooperation when their non-human identity is undisclosed. Yet, this advantage diminishes once their non-human status is revealed. The bias against agents persists and cannot be mitigated over time, even after humans realize that agents are more powerful in promoting human cooperation. Moreover, even when humans anticipate their counterparts to cooperate, they are more likely to exploit an agent’s benevolence than that of a human (Karpus et al., 2021). Such distinctions are also prevalent in emotional responses, particularly those related to feelings of guilt and envy towards agents and humans. Empirical results indicate that humans exhibit similar levels of envy when interacting with both agents and humans, but report significantly less guilt in interactions with agents (Melo et al., 2016).
Mitigating human bias.
Cultural cues, emotion, and verbal communication can reduce human biases in interactions with agents. de Melo and Terada (2019) organize human-agent experiments wherein humans participate in iterated PD with agents featuring virtual faces and emotional expressions. The results indicate that cultural and emotional features can lessen biases and promote more effective and cooperative interactions. Maggioni and Rossignoli (2023) further demonstrate that verbal communication in human-agent interactions in the iterated PD can encourage cooperative strategies, somewhat alleviating biases against agent counterparts. These findings underscore the importance of nuanced communication in fostering cooperation between humans and agents.
@@ 5 Inspirations on Human-Human Cooperation
In human-agent systems, agents scaffold human-human cooperation by assuming various roles. These include acting as planners to structure network interactions, operating as independent decision-makers to affect population composition, and making decisions on behalf of humans.
Engineering network structure.
@@
Social networks play a key role in overcoming SDs. Shirado and Christakis (2020) utilize intervention AI agents to reshape local social connections between humans. The results reveal a significant increase in human-human cooperation when agents employing disengaged intervention (cutting a defective neighbor) are introduced. Moreover, even a single agent implementing a mixed strategy (including engaged, disengaged, and self-rewiring intervention) significantly alters social dynamics and improves cooperation, fostering the development of cooperative clusters. Subsequently, McKee et al. (2023) explore human-human cooperation using deep reinforcement learning in network games. They train an agent as a GraphNet planner, responsible for guiding humans in creating or breaking interaction links. This planner consistently recommends building links between cooperators and discourages new links with defectors. Notably, the planner adopts a conciliatory strategy to address defectors, initially establishing a certain number of cooperate–defect links and progressively suggesting the severance of these links over time.
Influencing evolutionary dynamics.
@@
The evolution of cooperation is significantly influenced by the behavior, proportion, and spatial distribution of the agents. Terrucha et al. (2022) utilize evolutionary game theory to model the interactions and strategies of adaptive (human-like) and fixed-behavior agents in collective risk dilemma games. The findings suggest that the presence and behavior of agents significantly affect human cooperation levels, with humans adjusting their strategies to compensate for the agents’ actions. Motivated by this, Guo et al. (2023) examine the influence of agent composition on cooperation within human-agent systems. They find that a minority presence of agents can boost fully human-human cooperation. However, when agents form the majority, this may lead to a collapse in human cooperation. Additionally, agents with a purely constant strategy can effectively encourage human cooperation (e.g., loner agents (Sharma et al., 2023)). On the other hand, in situations where humans lack transparency regarding the actions of agents, Han et al. (2021) underscore the importance of considering the opportunity cost when adopting a reciprocal strategy. They develop a trust-based strategy that incorporates probabilistic checks to minimize potential opportunity costs, demonstrating its effectiveness in maintaining cooperation.
@@
Humans have the capacity to entrust decision-making authority to agents. Particularly, this mode of interaction can change the way humans solve these SDs. de Melo et al. (2019) find that humans tend to exhibit higher cooperation when decisions are agent-driven rather than made personally. The rationale is that programming encourages individuals to consider long-term interests while diminishing the influence of immediate short-term self-interest, thereby fostering greater cooperative behavior. Fernández Domingos et al. (2022) also find that cooperation and group success increase when participants delegate their actions to agents, allowing it to play on their behalf in a collective risk dilemma.
@@ 6 Directions of Future Research
(1) Enhancing cooperation in multi-agent and human-agent settings with large language models (LLMs), known for advanced reasoning, planning and abstracting capabilities (Park et al., 2023). LLMs can empower the study of cooperation in SDs in two ways. Firstly, they can serve as prior knowledge for shaping agent rewards, drawing inspiration from their recent use in (multi-agent) reinforcement learning (Hu and Sadigh, 2023). The second usage is to employ LLMs as the backbone of agent architectures for generative agents in SDs. Unlike existing work on fully cooperative settings using LLMs (Zhang et al., 2023a), creating generative agents that can achieve cooperation in SDs involves additional challenges, such as building reputation, identifying deception and avoiding exploitation from others, as well as nudging others to cooperate through potential punishment.
@@
(2) Establishing a theoretically guaranteed framework for fostering and sustaining cooperation in SDs. While methods surveyed in this paper show empirical promise in encouraging cooperation in SDs, none offers theoretical guarantees. A theoretical framework is envisioned to bridge past and future SD studies: precisely explaining existing empirical findings and guiding the design of more effective methods. It is crucial to note that conditions for maintaining cooperation differ between normal-form SDs and SSDs (Kleshnina et al., 2023), necessitating a separate consideration in constructing the theoretical framework. Moreover, for human-agent hybrid systems, existing studies are mostly empirical; those exceptions that leverage evolutionary game theory cannot capture human biases towards AI agents, a critical aspect of human-agent hybrid systems, and thus fail to provide an accurate prediction or explanation of the evolution of cooperation in these systems.
@@
(3) Further employing SDs to model real-world decision-making scenarios. Large-scale simulation systems on economics, sociology and ecology with diverse agents, e.g., the board game Diplomacy (Kramár et al., 2022), offer insights into the mechanism of system evolution. These low-cost, realistic simulations better inform policymakers and even inspire solutions to relief SDs. As AI agents are integrated with humans, they provide convenience whilst introduce possibly negative social influence to this human-agent hybrid system. A typical real-world scenario is autonomous driving, where AI agents face dilemmas like choosing between speed and safety at intersections and must navigate undesirable selfish behavior from human drivers. Addressing collision risks among autonomous and human-operated vehicles necessitates designing algorithms for large-scale human-agent hybrid systems. Such AI agents must adeptly respond to human behaviors and biases, underlining the importance of developing human-agent hybrid systems for realistic scenarios to enhance human-agent cooperation and tackle key cooperative challenges in SDs.
@@
(4) Bridging human-agent cooperation and SSDs. Compared to the fruitful literature on multi-agent cooperation in SSDs, the investigation of human-agent in SSDs is insufficient as of this survey, though with only few exceptions (Crandall et al., 2018). In SSDs, the process of decision-making is extended over a series of steps, consequently leading to human behavioral patterns that often diverge from those observed in the singe-stage settings (e.g., normal-form SDs) (Sims et al., 2013). This brings new challenges to human-agent cooperation: It will become difficult to develop AI systems that can nudge humans towards cooperation by avoiding being myopic and focusing on longer-term benefits, and to understand the behavioral tendencies humans exhibit in sequential decision-making processes.
(5) Revisiting existing theory of human cooperation, particularly in light of recent findings indicating that AI systems, such as autonomous safety features in vehicles, can potentially disrupt established norms of reciprocity among humans (Shirado et al., 2023). It has been shown that social norms are important mechanisms for maintaining human cooperation (Santos et al., 2018); however, these norms may be at risk of erosion when decision-making is potentially delegated to agents or when humans interact with agents. Such concerns extend to other mechanisms, such as network reciprocity and group identity. Thus, this calls for a rethinking of the existing theory of human cooperation, taking into account the growing integration of agents in human societies.

### v2

@@ 4 Human-Agent Cooperation in SDs
In the era of artificial intelligence, technologies like autonomous vehicles [81], chatbots [82], and human-computer interaction systems [83] have become ubiquitous. AI agents now handle delicate and complex tasks, ranging from disaster relief to city construction. While extensive research has explored how agents coordinate with humans - a field known as human-agent coordination [84, 85, 86] - there are also scenarios where conflicts of interest emerge between humans and agents. Human-agent coordination focuses on shared interests, where agents’ goals align perfectly with those of humans. In contrast, human-agent cooperation centers on exploring strategies to sustain high cooperation levels in human-agent hybrid systems amidst strategic conflicts and SDs. Despite overlaps between coordination and cooperation with agents, the primary challenges diverge. This section explores human-agent cooperation in SDs from two angles: (i) developing algorithms that effectively enhance cooperative behavior; (ii) the biases humans exhibit in their decision-making processes when interacting with both humans and agents. We summarize the references used in Section 4 at the top of Table 3.
@@ 4.1 Designing Algorithms for Cooperation with Humans
From S++ to S#
Successful AI agents require three key properties: generality across various games, adaptability to unfamiliar counterparts, and swift learning in decision-making [14]. The S++ and S# family addresses these needs, focusing on iterated interactions where individuals repeatedly engage in the same game structure and make decisions based on past experiences. [66, 67].
S++ [66], recognized as an expert algorithm, is originally designed for iterated normal-form games. It stands out with two main components: firstly, its set of experts, which are systematically generated from the game’s description, and secondly, its expert-selection mechanism, which dictates the choice of expert to be utilized in each round of the game. Building upon the S++ algorithm, Crandall et al. [14] systematically investigate human-agent cooperation in iterated normal-form games (e.g., PD and Chicken games) and stochastic games (e.g., Stochastic Game Prisoner’s Dilemma and Block Game). In the interaction with other algorithms such as generous Tit-for-Tat, Win-Stay-Lose-Shift, etc, the S++ algorithm emerges as the top performer. A human-agent experiment subsequently reveals that S++ is as proficient at establishing cooperative relationships with humans as humans are. However, S++ and humans often fail to achieve high cooperation levels with human partners. Therefore, a communication mechanism is introduced.
The S# algorithm [14], an advancement of the S++ algorithm, integrates a communication framework that allows agents to engage in and respond to cheap talk. Its capacity to enhance cooperation has been validated through both simulations and human-agent experiments. The results showcase its ability to cooperate with humans and other algorithms at levels that rival human cooperation. Subsequently, Oudah et al. [67] further develops the S++ and S# algorithms by introducing diverse signaling strategies inspired by various philosophical approaches. This amalgamation of behavior strategy (the agent’s actions) and signaling strategy (the agent’s verbal communication) endeavors to optimize player utility by influencing people and winning friends. Findings reveal that Carnegie’s Principles [87], a signaling strategy focused on refraining from criticism, complaints, or condemnation while positively uplifting partners, coupled with a behavior strategy characterized by rapid and effective learning, proves most adept at winning friends and influencing people.
@@
Mitigating the obstacle of cooperation requires handling human selfishness. Zero-determinant (ZD) strategy becomes a viable option. ZD strategies establish a linear payoff correlation between a focal player and the counterpart, irrespective of the counterpart’s strategy [88]. This category includes extortion strategies, ensuring that a player’s surplus consistently surpasses the counterpart’s surplus by a fixed percentage, and generosity strategies, designed to incentivize cooperation from the counterpart. Building upon extortion and generosity strategies, Hilbe et al. [15] conduct a human-agent experiment and observe that although extortion strategies do prevail over their human counterparts, this success is met with retaliation, leading humans to choose defection in response. Consequently, generosity emerges as the more profitable and effective strategy. In contrast, extortion strategies outperform generosity strategies when there is a long enough duration for interaction and when human participants are aware they are competing against a computerized counterpart [68]. Human-agent cooperation finds application in safety-critical interactions between humans and agents. Zhang et al. [69] find that conventional safe control methods may not consistently achieve the highest safety and performance when faced with self-seeking humans. Addressing the issue of myopic behaviors, they propose an algorithm incorporating the influence of risk and long-term reward. The algorithm can prevent autonomous agents from being exploited by self-seeking humans and seeks to achieve a more optimal balance between safety and performance in human-agent interactions.
Experiment-driven insights
Designing agents that promote cooperation can be inspired by experimental findings. Santos et al. [70] conduct human-agent experiments involving human participants who interact with two programmed agents in a collective risk dilemma. Following the game, participants complete questionnaires regarding their preferences for future robotic partners. The results from these post-task questionnaires indicate a tendency among humans to favor cooperating agents following collective failures. Conversely, there is no distinct preference following collective successes. This observation prompts the development of an outcome-based partner selection model, further explored through an evolutionary game theory framework to evaluate its effectiveness. An online human-agent study reinforces these findings and highlights the significant influence of outcome-based strategies in human-agent interactions [71].
@@ 4.2 Revealing Human Bias towards Agents
Bias caused by the nature of counterparts
Within human-agent systems, evidence has shown that the revelation of an agent’s true identity holds sway over human behavior [16, 72]. Ishowo-Oloko et al. [16] conduct an intriguing experiment involving iterated PD games, wherein participants or agents engage in continuous interactions with a designated counterpart. The crux of their investigation lies in probing the delicate tradeoff between transparency-revealing one’s identity as an agent—and efficacy-the capacity to cooperate seamlessly in human-agent interactions. Human participants, crucially, are divided into groups: one accurately briefed on the nature of their counterparts—whether human or agent and the other misled by false information. Employing the S++ algorithm, agents are more successful in eliciting human cooperation when their non-human identity is undisclosed. Yet, this advantage diminishes once their non-human status is revealed. The bias against agents persists and cannot be mitigated over time, even after humans realize that agents are more efficient in promoting human cooperation. Moreover, experiments involving SD games with human players and agent counterparts have shown that even when humans expect cooperation from their counterparts, they are more likely to exploit an agent’s benevolence than that of another human [72]. Notably, the agent’s behavior is engineered to mirror human actions in these games. Empirical findings reveal a noteworthy trend: participants don’t demonstrate diminished trust in agents relative to humans. Instead, they display a heightened tendency to exploit agents, particularly when anticipating cooperation from the opposing party. Such distinctions are also prevalent in emotional responses, particularly those related to feelings of guilt and envy towards agents and humans. Melo et al. [73] reveal that humans exhibit similar levels of envy when interacting with both agents and humans, but report significantly less guilt in interactions with agents within SD games.
Mitigating human bias
Cultural cues, emotion, and verbal communication hold promise for mitigating human biases in interactions with agents. de Melo and Terada [17] organize human-agent experiments wherein humans participate in iterated PD with agents featuring virtual faces and emotional expressions. The findings unveiled a compelling insight: cultural and emotional attributes have the potential to diminish biases, fostering more effective and cooperative interactions. Subsequently, Maggioni and Rossignoli [74] delve deeper into the realm of human-agent interaction within iterated PD games, demonstrating the influential role of verbal communication in fostering cooperative strategies and partially mitigating biases against agent counterparts. Together, these revelations offer a glimmer of hope in mitigating biases against agent counterparts.
@@ 5 Inspirations on Human-Human Cooperation
In human-agent systems, agents scaffold human-human cooperation by assuming various roles. These include acting as planners to structure network interactions, operating as independent decision-makers to affect population composition, and making decisions on behalf of humans. We summarize the references used in Section 5 at the bottom of Table 3.
Engineering network structure
@@
Social networks play a key role in overcoming SDs. Shirado and Christakis [75] utilize intervention AI agents to reshape local social connections between humans. The results reveal a significant increase in human-human cooperation when agents employing disengaged intervention (cutting a defective neighbor) are introduced. Moreover, even a single agent implementing a mixed strategy (including engaged, disengaged, and self-rewiring intervention) significantly alters social dynamics and improves cooperation, fostering the development of cooperative clusters. Subsequently, McKee et al. [18] explore human-human cooperation using deep reinforcement learning in network games. They train an agent as a GraphNet planner, responsible for guiding humans in creating or breaking interaction links. This planner consistently recommends building links between cooperators and discourages new links with defectors. Notably, the planner adopts a conciliatory strategy to address defectors, initially establishing a certain number of cooperate–defect links and progressively suggesting the severance of these links over time.
Influencing evolutionary dynamics
@@
In human-agent populations, the evolution of cooperation hinges significantly on the behavior, proportion, and spatial distribution of the agents. Terrucha et al. [76] utilize evolutionary game theory to model the interactions and strategies of adaptive (human-like) and fixed-behavior agents in collective risk dilemma games. Their findings suggest that the presence and behavior of agents significantly affect human cooperation levels, with humans adjusting their strategies to compensate for the agents’ actions. Motivated by this, Guo et al. [19] investigate the impact of cooperative and defective agents on human cooperation. Employing replicator dynamics and pairwise comparison, the study shows cooperative agents have varying impacts, with limited influence in prisoner’s dilemma games but promoting cooperation in stag hunt games. Defective agents, intriguingly, can lead to complete dominance of cooperation in snowdrift games. Additionally, the role of population structure and imitation strength is highlighted as crucial in these dynamics. Agents with a purely constant strategy can also efficiently encourage human cooperation. Sharma et al. [77] discover that certain types of agents, such as always-cooperate agents and loner agents, significantly promote cooperation under different structural conditions. In particular, this study identifies an optimal density of loner agents for maximizing cooperation, suggesting that the diversity in agent actions can diminish their cooperation-promoting effect. On the other hand, in situations where humans lack transparency regarding the actions of agents, Han et al. [78] underscore the importance of considering the opportunity cost when adopting a reciprocal strategy. In response to this concern, they develop a trust-based strategy that, upon establishing trust, utilizes probabilistic checks to minimize potential opportunity costs, demonstrating its effectiveness in maintaining cooperation.
@@
Humans have the capacity to entrust decision-making authority to agents. Particularly, this mode of interaction can change the way humans solve these SDs. de Melo et al. [79] find that humans exhibit higher cooperation when decisions are agent-driven rather than made personally in games with repeated interactions. The rationale is that programming encourages individuals to consider long-term interests while diminishing the influence of immediate short-term self-interest, thereby fostering greater cooperative behavior. Fernández Domingos et al. [80] also highlight the positive impact of delegating decision-making to AI agents in collective risk scenarios, including two forms of delegation: to AI agents with predefined behaviors and to customizable AI agents whose behaviors are defined by participants. Both forms significantly enhanced cooperation and group success in achieving collective targets, compared to groups composed solely of humans. Notably, customizable AI agents led participants to adopt cooperative strategies more frequently, underscoring the potential of AI as a tool for promoting prosocial behavior.
@@ 6 Directions of Future Research
(1) Enhancing cooperation in multi-agent and human-agent settings with large language models (LLMs). LLMs are generative AI models trained on extensive human textual data [89]. LLMs excel in processing and generating natural language text, which makes them useful for a wide range of language tasks. Therefore, LLMs have been employed to construct large social simulation systems with higher interpretability [90, 91]. When focusing on the study of cooperation in SDs, LLMs might empower the study in the following two ways. Firstly, they can serve as a source of prior knowledge for shaping agent rewards. LLMs have been employed to generate rewards or reward functions for RL agents in recent works, steering RL agents to coordinate with humans better or achieve human-level precision control [92, 93, 94, 95]. Nevertheless, it remains an open question how RL agents that learn from LLM-shaped rewards to solve SDs effectively. The second usage is to employ LLMs as the backbone of agent architectures for generative agents in SDs. Some recent works have evaluated the traits and behavioral characteristics of LLMs in SD games (e.g. PD, SH and HD), but their conclusions reached are not always consistent [96, 97, 98, 99, 100]. The differences in their conclusions may be due to variations in the prompts they used. Continuing to explore the capabilities and performance of LLMs in SDs is crucial, especially as we tackle additional challenges in these issues like building reputation, identifying deception, and avoiding exploitation, as well as encouraging cooperation through potential punitive measures.
@@
(2) Establishing a theoretically guaranteed framework for fostering and sustaining cooperation in SDs. While methods surveyed in this paper show empirical promise in encouraging cooperation in SDs, none offers theoretical guarantees. A theoretical framework is envisioned to bridge past and future SD studies: precisely explaining existing empirical findings and guiding the design of more effective methods. It is crucial to note that conditions for maintaining cooperation differ between normal-form SDs and SSDs [101], necessitating a separate consideration in constructing the theoretical framework. Moreover, for human-agent hybrid systems, existing studies are mostly empirical those exceptions that leverage evolutionary game theory cannot capture human biases towards AI agents, a critical aspect of human-agent hybrid systems, and thus fail to provide an accurate prediction or explanation of the evolution of cooperation in these systems. It is worth emphasizing that when handling behavioral experiment data from human users, preserving user privacy through theoretical methods is also crucial[102, 103, 104].
@@
(3) Further employing SDs to model real-world decision-making scenarios. Large-scale simulation systems on economics, sociology and ecology with diverse agents, e.g., the board game Diplomacy [105], offer insights into the mechanism of system evolution. These low-cost, realistic simulations better inform policymakers and even inspire solutions to relief SDs. As AI agents are integrated with humans, they provide convenience whilst introduce possibly negative social influence to this human-agent hybrid system. A typical real-world scenario is autonomous driving, where AI agents face dilemmas like choosing between speed and safety at intersections and must navigate undesirable selfish behavior from human drivers. Current SD research often simplifies driving scenarios [36]. As autonomous vehicles being integrated with humans, infusing them with realistic social characteristics is crucial for positive societal impact. Addressing collision risks among autonomous and human-operated vehicles necessitates designing algorithms for large-scale human-agent hybrid systems and constructing cost-effective, extensive simulation systems. More broadly, SDs holds promise to build simulation systems spanning economics, sociology, or ecology can simulate AI agents with diverse social preferences, offering insights into system evolution. These realistic simulations can better inform policymakers. However, it is important to note that interpretations of AI simulation outcomes requires a dialectical perspective, and vigilance toward ethical concerns is paramount. All in all, such SDs-based AI agents must adeptly respond to human behaviors and biases, underlining the importance of developing human-agent hybrid systems for realistic scenarios to enhance human-agent cooperation and tackle key cooperative challenges in SDs.
@@
(4) Bridging human-agent cooperation and SSDs. Compared to the fruitful literature on multi-agent cooperation in SSDs, the investigation of human-agent in SSDs is insufficient as of this survey, though with only few exceptions [14]. In SSDs, the process of decision-making is extended over a series of steps, consequently leading to human behavioral patterns that often diverge from those observed in the single-stage settings (e.g., normal-form SDs) [106]. This brings new challenges to human-agent cooperation: It will become difficult to develop AI systems that can nudge humans towards cooperation by avoiding being myopic and focusing on longer-term benefits, and to understand the behavioral tendencies humans exhibit in sequential decision-making processes. In SSD scenarios, people must consider both their behavior rules, the behavior of agents, and their interactions with the environment, which form a complex feedback loop. For example, cooperative behaviors can cultivate a peaceful or replete environment, leading to long-term benefits. However, the need to maximize immediate profits adds complexity to decision-making. This complexity is further driven by the dynamic nature of the environment, which continuously interacts with and shapes human behaviors. Understanding this interplay is essential for devising effective cooperation strategies in SSDs, highlighting the need to balance behavior rules and immediate gains with long-term considerations.
(5) Revisiting existing theory of human cooperation, particularly in light of recent findings indicating that AI systems, such as autonomous safety features in vehicles, can potentially disrupt established norms of reciprocity among humans [107]. Traditionally, theories such as social norms [108], prosocial preferences [109], and reciprocity mechanisms [1] have been pivotal in explaining cooperative behaviors among humans and other species. However, the emergence of a decline in reciprocity observed in semi-automatic driving [107] prompts a deeper inquiry into the applicability of these theories within contexts where humans interact with agents. For instance, prosocial preferences suggest that individuals cooperate because they value the welfare of others; yet, it remains an open question whether such preferences extend to interactions involving non-human agents [110, 111, 112]. Will individuals exhibit prosocial behaviors when their counterparts are perceived not as fellow humans but as agents? Can prosocial preferences still promote cooperation in environments where humans interact with AI? Similar inquiries are pertinent for other cooperative mechanisms, such as network reciprocity and both direct and indirect forms of reciprocity [5]. This evolving landscape calls for a critical reevaluation of the traditional theories of human cooperation, especially considering the growing integration of AI agents into human societies.
Acknowledgement
This work was supported in part by the National Science Fund for the National Key R&D Program (No. 2022YFE0112300), Distinguished Young Scholarship of China (No. 62025602), the National Natural Science Foundation of China (Nos. U22B2036, 11931015), Fok Ying-Tong Education Foundation China (No. 171105), the Fundamental Research Funds for the Central Universities (No. G2024WD0151), the Tencent Foundation and XPLORER PRIZE, China Postdoctoral Science Foundation (No. 2023M741852), JSPS Postdoctoral Fellowship Program for Foreign Researchers (No. P21374), JSPS KAKENHI (No. JP 22KF0303), and Shanghai Artificial Intelligence Laboratory.